Find the exact length of the side of a square whose diagonal is 3 feet.
Answer by RAY100(1637)

The diagonal of the square makes a triangle that is ,,45 - 45 - 90 degrees,
with sides,,,,,1-1-sqrt2
.
sqrt2 / 3 = 1/x
.
cross multiply
.
xsqrt2 =3
.
x=3/sqrt2 = 3sqrt2 / 2 = 2.12,,,,,,,therefore side = (3/2) sqrt2
.
check
sqrt(2.12^2 +2.12^2) = 3,,,,,ok